1. Calculate the turning points of the function y=sin3t using differential calculus
  2. Show which are maxima, minima or points of inflexion using the second derivative
1
Expert's answer
2021-11-09T00:01:19-0500

1. "y=\sin(3t)"

Domain: "(-\infin, \infin)"

Find the first derivative


"y'=(\sin(3t))'=3\cos(3t)"

Find the critical number(s)


"y'=0=>3\cos(3t)=0"

"3t=\dfrac{\pi}{2}+\pi n, n\in \Z"

"t=\dfrac{\pi}{6}+\dfrac{\pi n}{3}, n\in \Z"

"y(\dfrac{\pi}{6}+\dfrac{2\pi m}{3}, m\in \Z)=1"

"y(\dfrac{\pi}{2}+\dfrac{2\pi k}{3}, k\in \Z)=-1"

The turning points are

"\big(\dfrac{\pi}{6}+\dfrac{2\pi m}{3}, 1\big), \big(\dfrac{\pi}{2}+\dfrac{2\pi k}{3},-1\big), m,k\in \Z"

2.

Find the second derivative


"y''=(3\cos(3t))'=-9\sin(3t)"

"y''(\dfrac{\pi}{6}+\dfrac{2\pi m}{3})=-9<0"

"y''(\dfrac{\pi}{2}+\dfrac{2\pi m}{3})=9>0"

The points "\big(\dfrac{\pi}{6}+\dfrac{2\pi m}{3}, 1\big), m\in \Z" are maxima.


The points "\big(\dfrac{\pi}{2}+\dfrac{2\pi k}{3}, 1\big), k\in \Z" are minima.



"y''=0=>-9\sin(3t)=0"

"3t=\pi l, l\in \Z"

"t=\dfrac{\pi l}{3}, l\in \Z"

"y(\dfrac{\pi l}{3})=0, l\in \Z"

The points "\big(\dfrac{\pi l}{3}, 0\big), l\in \Z" are points of inflection.